# Regularity Properties of Optimization-Based Controllers

### v4

@@ 3 Regularity of Parametric Optimizers
Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strongly convex. Suppose that SC holds at $x_{0}\in^{n}$. Then, $u^{*}$ is locally Lipschitz at $x_{0}$.
@@
Note that for all $i\in\mathcal{I}(x_{0},u^{*}(x_{0}))$, $g_{i}^{0}(x_{0})\neq 0$. Indeed, if $g_{i}^{0}(x_{0})=0$ and $i\in\mathcal{I}(x_{0},u^{*}(x_{0}))$, it follows that $g_{i}^{1}(x_{0})=0$, which implies that Slater’s condition at $x_{0}$ is violated. Hence, $g_{i}^{0}(x_{0})\neq 0$ for all $i\in\mathcal{I}(x_{0},u^{*}(x_{0}))$ and the CR holds at $(x_{0},u^{*}(x_{0}))$. Moreover, since Slater’s condition holds at $x_{0}$, by [59, Prop. 5.39], since $g(x_{0},\cdot)$ is convex, MFCQ holds at $(x_{0},u^{*}(x_{0}))$. By [55, Theorem 3.6], this implies that $u^{*}$ is locally Lipschitz at $x_{0}$. ∎
Note that Robinson’s counterexample or [55, Example 3.11] do not contradict Proposition 3.3, since in those two examples the optimization variable of the quadratic program has dimensions four and two, respectively. Proposition 3.3 shows that optimization-based controllers for single-input systems with affine constraints (e.g., obtained from CBF or CLF based conditions for control-affine systems) that satisfy Slater’s conditions are locally Lipschitz.
@@
(Not point-Lipschitz optimizer without differentiability of problem data with respect to the parameter): If $f$ and $g$ are not differentiable with respect to the parameter $x$ but the rest of the assumptions of Proposition 3.2 hold, the following example, inspired by Robinson’s counterexample, shows that the parametric optimizer is not necessarily point-Lipschitz. Let $x=(x_{1},x_{2})\in^{2}$ and consider (22) with
@@
Since $x_{1}$ can be taken to be arbitrarily small, $\tilde{u}^{*}$ is not point-Lipschitz at the origin. However, because $f$ and $g$, as well as their first and second derivatives with respect to $u$, are continuous in $u$ and $x$, and the rest of assumptions of Proposition 3.2 hold, then by [64, Theorem 5.3], the corresponding parametric optimizer, and hence $\tilde{u}_{4}^{*}$, is continuous. $\bullet$
@@
(Discontinuous optimizer without Slater’s condition): The following example, taken from [42, Section VI], shows that if Slater’s condition does not hold, then continuity of the parametric optimizer is not guaranteed even if the rest of assumptions from Proposition 3.2 do hold:
@@
(Unbounded optimizer without Slater’s condition): The following example, adapted from [48, Example III.5], shows that if Slater’s condition fails, not only can the parametric optimizer fail to be continuous, as shown in Example 3.5, but it can even fail to be locally bounded. Let $x=(x_{1},x_{2})\in^{2}$, $a(x)=2x_{1}x_{2}+x_{2}^{2}(1-x_{1}^{2}-x_{2}^{2})$, and consider:
@@
Note that since $a(1,0)=0$, any neighborhood of $(1,0)$ contains a point $x_{+}$ such that $a(x_{+})>0$ and a point $x_{-}$ such that $a(x_{-})\leq 0$. Moreover, since
$\lim\limits_{(x_{1},x_{2})\to(1,0)}-\frac{a(x)}{2x_{2}^{3}}=\infty,$
it follows that $\breve{u}^{*}$ is not locally bounded. $\bullet$
@@
(Conditions for local boundedness): Suppose that $f$ and $g$ belong to $\mathcal{C}^{2,2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strongly convex, $g(x,\cdot)$ is convex, and that the feasible set $\left\{u\in^{m}\;\mid\;g(x,u)\leq 0\right\}$ is nonempty. Then, given $x_{0}\in^{n}$, $u^{*}$ is locally bounded at $x_{0}$ if and only if LCF holds at $x_{0}$.
@@
First suppose that LCF holds at $x_{0}$. Therefore, there exists a compact set $K\subset^{m}$ and $\delta>0$ such that for all $y\in^{n}$ such that $\left\lVert y-x\right\rVert<\delta$, there exists $u\in K$ such that $g(y,u)\leq 0$. Since $f$ is continuous and $K$ is compact, there exists $B_{f}>0$ such that $|f(y,u)|<B_{f}$ for all $u\in K$ and $y\in^{n}$ such that $\left\lVert y-x_{0}\right\rVert<\delta$. Since for all $y\in^{n}$ such that $\left\lVert y-x_{0}\right\rVert<\delta$, there exists a feasible $u\in K$, it follows that $|f(y,u^{*}(y))|<B_{f}$ for all $y\in^{n}$ such that $\left\lVert y-x_{0}\right\rVert<\delta$. This implies that $u^{*}$ is locally bounded at $x_{0}$. Now suppose that $u^{*}$ is locally bounded at $x_{0}$ and suppose, by contradiction, that LFC does not hold at $x_{0}$. Then, for any $\delta>0$ and compact set $K$, there exists $y\in^{n}$ with $\left\lVert y-x\right\rVert\leq\delta$ and such that all $u\in^{m}$ with $g(u,y)\leq 0$ satisfy $u\notin K$. This means that there exists a sequence $\{y_{n}\}_{n\in\mathbb{Z}_{>0}}$ such that $\left\lVert y_{n}-x\right\rVert\leq 1/n$ and $\left\lVert u^{*}(y_{n})\right\rVert\geq n$ for all $n\in\mathbb{Z}_{>0}$, which implies that $u^{*}$ is not locally bounded, hence reaching a contradiction. ∎
Verifying the local compact feasibility property can be challenging in general. However, for the particular case of CBF-based quadratic programs, [48, Theorem V.1] gives an alternative sufficient condition for local boundedness of $u^{*}$ that only requires solving a specific linear equation.
@@
(Sufficient conditions for measurability): Suppose that $f$ and $g$ belong to $\mathcal{C}^{2,2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strongly convex, $g(x,\cdot)$ is convex, and that the feasible set $\left\{u\in^{m}\;\mid\;g(x,u)\leq 0\right\}$ is nonempty. Further assume that for every $x\in^{n}$ LCF holds at $x$. Then, $u^{*}$ is measurable.
@@
We use the Measurable Maximum Theorem [65, Theorem 18.19]. We assume that n and m are equipped with the usual Borel $\sigma$-algebras. Since $f$ belongs to $\mathcal{C}^{2,2}(^{n}\times^{m})$, it is a Carathéodory function (cf. [65, Definition 4.50]). Therefore, we only need to ensure that the set-valued map $\phi:x\to\left\{u\in^{m}\;\mid\;g(u,x)\leq 0\right\}$ is a weakly measurable correspondence (cf. [65, Definition 18.1]) with nonempty compact values. The fact that $\phi$ takes nonempty values follows from the fact that the feasible set $\left\{u\in^{m}\;\mid\;g(x,u)\leq 0\right\}$ is nonempty. Moreover, since Proposition 3.7 ensures that $u^{*}$ is locally bounded at every $x\in^{n}$, without loss of generality we can assume that $\phi$ takes compact values (otherwise, we can define extra constraints that ensure that the feasible set is bounded for every $x\in^{n}$ without changing the optimizer $u^{*}$). Now, to show that $\phi$ is a weakly measurable correspondence, we follow an argument similar to the proof of [65, Corollary 18.8]. For every $n\in\mathbb{Z}_{>0}$, define the set-valued map $\phi_{n}:x\to\left\{u\in^{m}\;\mid\;g(u,x)\leq 1/n\right\}$. By Lemma [65, Corollary 18.8], $\phi_{n}$ is measurable. Moreover, for every $x\in^{n}$ and $n\in\mathbb{Z}_{>0}$, $\phi(x)\subset\partial(\phi_{n}(x))$ (where $\partial(\phi_{n}(x))$ denotes the boundary of $\phi_{n}(x)$), and $\phi(x)=\cap_{n=1}^{\infty}\partial(\phi_{n}(x))$. Furthermore, again without loss of generality, $\partial(\phi_{n})$ has compact values for every $n\in\mathbb{Z}_{>0}$ (again, if that is not the case we can define extra constraints that ensure that this holds without changing the optimizer $u^{*}$), and by [65, Theorem 18.4(3)], the intersection $\phi:x\to\cap_{n=1}^{\infty}\partial(\phi_{n}(x))$ is measurable. ∎
@@
(Verifying constraint qualifications and conditions in practice without knowledge of the optimizer): To show that $u^{*}$ is locally Lipschitz at a point $x\in^{n}$ using  [55, Theorem 3.6], we need to verify that both MFCQ and CR hold at $(x,u^{*}(x))$. Similarly, [54, Theorem 4.1] (resp. [53, Theorem 2.1]) require the verification of LICQ (resp. LICQ and SCS) at $(x,u^{*}(x))$. These results require knowledge of $u^{*}(x)$ to verify the corresponding property holds at $x$. However, in several applications it can be useful to know the regions where the controller $u^{*}$ is discontinuous (for instance, to design safety-critical controllers that avoid such regions). Slater’s condition is useful for this purpose because Proposition 3.2 guarantees different regularity properties of $u^{*}$ at $x$ without requiring knowledge of $u^{*}(x)$ (assuming that the extra conditions on differentiability and convexity of the objective function and constraints in Proposition 3.2 also hold). Moreover, in the special case where the constraints in (1) are affine, i.e.,
@@ 4 Existence and Uniqueness of Solutions under Optimization-Based Controllers
(Existence of classical solutions for the closed-loop system):  Suppose that $f$ and $g$ belong to $\mathcal{C}^{2,2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(\cdot,x)$ is strongly convex, $g(\cdot,x)$ is convex and SC holds at $x_{0}\in^{n}$. Let $F:^{n}\times^{m}\to^{n}$ be locally Lipschitz. Then, there exists $\delta_{0}>0$ such that the differential equation (18) has at least one solution $x:(-\delta_{0},\delta_{0})\to^{n}$ with initial condition $x(0)=x_{0}$.
@@
By Proposition 3.2, $u^{*}$ has the Hölder property at $x_{0}$ and there exists a neighborhood $\mathcal{V}_{x_{0}}$ of $x_{0}$ such that $u^{*}$ is point-Lipschitz at $y$ for all $y\in\mathcal{V}_{x_{0}}$. Both of these properties imply that $u^{*}$ is continuous in a neighborhood of $x_{0}$. The result follows by Peano’s existence theorem [67, Theorem 2.1]. ∎
@@
Figure 2: The blue arrows depict the vector field (26). The dashed red and green curves depict the two solutions $y_{1}$ and $y_{2}$ starting from the origin, where the vector field is point-Lipschitz but not locally Lipschitz.
@@
If $\tilde{F}(x_{0})\neq 0$, by [68, Theorem 1.2.7], the differential equation has only one solution. If $\tilde{F}(x_{0})=0$, the result follows from Proposition 4.3. ∎
If Slater’s condition does not hold but the rest of assumptions of Proposition 4.1 hold, Example 3.5 shows that $u^{*}$ can be discontinuous, in which case neither existence nor uniqueness of solutions is guaranteed. In the case where $f$ and $g$ are not differentiable with respect to the parameter, but the rest of assumptions of Proposition 4.1 hold, Example 3.4 shows that $u^{*}$ is continuous but not necessarily point-Lipschitz. Therefore, in this case existence is guaranteed but uniqueness is not.
@@
(Existence of Filippov solutions for the closed-loop system): Suppose that $f$ and $g$ belong to $\mathcal{C}^{2,2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(\cdot,x)$ is strongly convex, $g(\cdot,x)$ is convex and the feasible set is nonempty. Finally suppose that for every $x\in^{n}$, LCF holds at $x$ and $F:^{n}\times^{m}\to^{n}$ is locally Lipschitz. Then for any $x\in^{n}$, there exists $\delta_{x}>0$ such that (18) has at least one Filippov solution $y:(-\delta_{x},\delta_{x})\to^{n}$ with initial condition $y(0)=x$.
@@
By Propositions 3.7 and 3.8, the assumptions ensure that $u^{*}$ is measurable and locally bounded. The result follows from [69, Theorem 7]. ∎
@@
Finally, we note that if $f$ is not strongly convex or $g$ is not convex, the optimizer $u^{*}$ is not guaranteed to be single-valued, which means that the usual notions of regularity of the controller and of solutions of the closed-loop system are not well defined. The third and fourth columns of Table 1 summarize the results presented in this section.

### v5

@@ 3 Regularity of Parametric Optimizers
Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strictly convex and (1) has at least one minimizer. Suppose that SC holds at $x_{0}\in^{n}$. Then, $u^{*}$ is locally Lipschitz at $x_{0}$.
@@
First, since (1) has at least one minimizer for all $x\in^{n}$, the convexity assumptions of $f$ and $g$ imply that $u^{*}$ is a singleton for all $x\in^{n}$. Note that for all $i\in\mathcal{I}(x_{0},u^{*}(x_{0}))$, $g_{i}^{0}(x_{0})\neq 0$. Indeed, if $g_{i}^{0}(x_{0})=0$ and $i\in\mathcal{I}(x_{0},u^{*}(x_{0}))$, it follows that $g_{i}^{1}(x_{0})=0$, which implies that Slater’s condition at $x_{0}$ is violated. Hence, $g_{i}^{0}(x_{0})\neq 0$ for all $i\in\mathcal{I}(x_{0},u^{*}(x_{0}))$ and the CR holds at $(x_{0},u^{*}(x_{0}))$. Moreover, since Slater’s condition holds at $x_{0}$, by [62, Prop. 5.39], since $g(x_{0},\cdot)$ is convex, MFCQ holds at $(x_{0},u^{*}(x_{0}))$. By [54, Theorem 3.6], this implies that $u^{*}$ is locally Lipschitz at $x_{0}$. ∎
Note that Robinson’s counterexample or [54, Example 3.11] do not contradict Proposition 3.3, since in those two examples the optimization variable of the quadratic program has dimensions four and two, respectively. Proposition 3.3 shows that optimization-based controllers for single-input systems with affine constraints (e.g., obtained from CBF or CLF based conditions for control-affine systems) that satisfy Slater’s conditions are locally Lipschitz.
@@
(Not point-Lipschitz optimizer without differentiability of problem data with respect to the parameter): If $f$ and $g$ are not differentiable with respect to the parameter $x$ but the rest of the assumptions of Proposition 3.2 hold (even with $f(x,\cdot)$ strongly convex for all $x\in^{n}$), the following example, inspired by Robinson’s counterexample, shows that the parametric optimizer is not necessarily point-Lipschitz. Let $x=(x_{1},x_{2})\in^{2}$ and consider (22) with
@@
Since $x_{1}$ can be taken to be arbitrarily small, $\tilde{u}^{*}$ is not point-Lipschitz at the origin. However, because $f$ and $g$, as well as their first and second derivatives with respect to $u$, are continuous in $u$ and $x$, and the rest of assumptions of Proposition 3.2 hold (with $f(x,\cdot)$ strongly convex for all $x\in^{n}$), then by [55, Theorem 5.3], the corresponding parametric optimizer, and hence $\tilde{u}_{4}^{*}$, is continuous. $\bullet$
@@
(Discontinuous optimizer without Slater’s condition): The following example, taken from [41, Section VI], shows that if Slater’s condition does not hold, then continuity of the parametric optimizer is not guaranteed even if the rest of assumptions from Proposition 3.2 (even with $f(x,\cdot)$ strongly convex for all $x\in^{n}$) do hold:
@@
(Unbounded optimizer without Slater’s condition): The following example, adapted from [47, Example III.5], shows that if Slater’s condition fails, not only can the parametric optimizer fail to be continuous, as shown in Example 3.5, but it can even fail to be locally bounded. Let $x=(x_{1},x_{2})\in^{2}$, $a(x)=2x_{1}x_{2}+x_{2}^{2}(1-x_{1}^{2}-x_{2}^{2})$, and consider:
@@
Note that $a(1,0)=0$ and $a(1,\epsilon)=2\epsilon-\epsilon^{4}$. Moreover, since any neighborhood of $(1,0)$ contains points of the form $(1,\epsilon)$ for sufficiently small $\epsilon>0$, for any neighborhood $\mathcal{N}$ of $(1,0)$ there exists $\epsilon_{\mathcal{N}}>0$ sufficiently small such that $a(1,\epsilon_{\mathcal{N}})>0$. Now, since
$\lim\limits_{\epsilon\to 0}\frac{a(1,\epsilon)}{2\epsilon^{3}}=\infty,$
and $\breve{u}^{*}(x)=-\frac{a(x)}{2x_{2}^{3}}$ if $a(x)>0$, it follows that $\breve{u}^{*}$ is not locally bounded. $\bullet$
@@
(Conditions for local boundedness): Suppose that $f$ and $g$ belong to $\mathcal{C}^{0}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strictly convex, $g(x,\cdot)$ is convex, and (1) has at least one minimizer. Then, given $x_{0}\in^{n}$, $u^{*}$ is locally bounded at $x_{0}$ if and only if LCF holds at $x_{0}$.
@@
Note that since (1) has at least one minimizer for all $x\in^{n}$, the convexity assumptions on $f$ and $g$ imply that $u^{*}$ is a singleton for all $x\in^{n}$. First suppose that LCF holds at $x_{0}$. Therefore, there exists a compact set $K\subset^{m}$ and $\delta>0$ such that for all $y\in^{n}$ such that $\left\lVert y-x\right\rVert<\delta$, there exists $u\in K$ such that $g(y,u)\leq 0$. Since $f$ is continuous and $K$ is compact, there exists $B_{f}>0$ such that $|f(y,u)|<B_{f}$ for all $u\in K$ and $y\in^{n}$ such that $\left\lVert y-x_{0}\right\rVert<\delta$. Since for all $y\in^{n}$ such that $\left\lVert y-x_{0}\right\rVert<\delta$, there exists a feasible $u\in K$, it follows that $|f(y,u^{*}(y))|<B_{f}$ for all $y\in^{n}$ such that $\left\lVert y-x_{0}\right\rVert<\delta$. This implies that $u^{*}$ is locally bounded at $x_{0}$. Now suppose that $u^{*}$ is locally bounded at $x_{0}$ and suppose, by contradiction, that LFC does not hold at $x_{0}$. Then, for any $\delta>0$ and compact set $K$, there exists $y\in^{n}$ with $\left\lVert y-x\right\rVert\leq\delta$ and such that all $u\in^{m}$ with $g(y,u)\leq 0$ satisfy $u\notin K$. This means that there exists a sequence $\{y_{n}\}_{n\in\mathbb{Z}_{>0}}$ such that $\left\lVert y_{n}-x\right\rVert\leq 1/n$ and $\left\lVert u^{*}(y_{n})\right\rVert\geq n$ for all $n\in\mathbb{Z}_{>0}$, which implies that $u^{*}$ is not locally bounded, hence reaching a contradiction. ∎
Verifying the local compact feasibility property can be challenging in general. However, for the particular case of CBF-based quadratic programs, [47, Theorem V.1] gives an alternative sufficient condition for local boundedness of $u^{*}$ that only requires solving a specific linear equation.
@@
(Sufficient conditions for measurability): Suppose that $f$ and $g$ belong to $\mathcal{C}^{0}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strictly convex, $g(x,\cdot)$ is convex, and (1) has at least one minimizer. Further assume that for every $x\in^{n}$, LCF holds at $x$. Then, $u^{*}$ is measurable.
@@
Note that since (1) has at least one minimizer for all $x\in^{n}$, the convexity assumptions on $f$ and $g$ imply that $u^{*}$ is a singleton for all $x\in^{n}$. We use the Measurable Maximum Theorem [67, Theorem 18.19]. We assume that n and m are equipped with the usual Borel $\sigma$-algebras. Since $f$ is continuous, it is a Carathéodory function (cf. [67, Definition 4.50]). Therefore, we only need to ensure that the set-valued map $\phi:x\to\left\{u\in^{m}\;\mid\;g(x,u)\leq 0\right\}$ is a weakly measurable correspondence (cf. [67, Definition 18.1]) with nonempty compact values. The fact that $\phi$ takes nonempty values follows from the fact that the feasible set $\left\{u\in^{m}\;\mid\;g(x,u)\leq 0\right\}$ is nonempty. Moreover, since Proposition 3.7 ensures that $u^{*}$ is locally bounded at every $x\in^{n}$, without loss of generality we can assume that $\phi$ takes compact values (otherwise, we can define extra constraints that ensure that the feasible set is bounded for every $x\in^{n}$ without changing the optimizer $u^{*}$). Now, to show that $\phi$ is a weakly measurable correspondence, we follow an argument similar to the proof of [67, Corollary 18.8]. For every $n\in\mathbb{Z}_{>0}$, define the set-valued map $\phi_{n}:x\to\left\{u\in^{m}\;\mid\;g(x,u)\leq 1/n\right\}$. By Lemma [67, Corollary 18.8], $\phi_{n}$ is measurable. Moreover, for every $x\in^{n}$ and $n\in\mathbb{Z}_{>0}$, $\phi(x)\subset\partial(\phi_{n}(x))$ (where $\partial(\phi_{n}(x))$ denotes the boundary of $\phi_{n}(x)$), and $\phi(x)=\cap_{n=1}^{\infty}\partial(\phi_{n}(x))$. Furthermore, again without loss of generality, $\partial(\phi_{n})$ has compact values for every $n\in\mathbb{Z}_{>0}$ (again, if that is not the case we can define extra constraints that ensure that this holds without changing the optimizer $u^{*}$), and by [67, Theorem 18.4(3)], the intersection $\phi:x\to\cap_{n=1}^{\infty}\partial(\phi_{n}(x))$ is measurable. ∎
@@
(Verifying constraint qualifications and conditions in practice without knowledge of the optimizer): To show that $u^{*}$ is locally Lipschitz at a point $x\in^{n}$ using  [54, Theorem 3.6], we need to verify that both MFCQ and CR hold at $(x,u^{*}(x))$. Similarly, [53, Theorem 4.1] (resp. [52, Theorem 2.1]) require the verification of LICQ (resp. LICQ and SCS) at $(x,u^{*}(x))$. These results require knowledge of $u^{*}(x)$ to verify the corresponding property holds at $x$. However, in several applications it can be useful to know the regions where the controller $u^{*}$ is discontinuous (for instance, to design safety-critical controllers that avoid such regions). Slater’s condition is useful for this purpose because Proposition 3.2 guarantees different regularity properties of $u^{*}$ at $x$ without requiring knowledge of $u^{*}(x)$ (assuming that the extra conditions on differentiability and convexity of the objective function and constraints in Proposition 3.2 also hold). Moreover, in the special case where the constraints in (1) are affine, i.e.,
@@ 4 Existence and Uniqueness of Solutions under Optimization-Based Controllers
(Existence of classical solutions for the closed-loop system):  Suppose that $f$ and $g$ belong to $\mathcal{C}^{0,2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strictly convex, $g(x,\cdot)$ is convex, and (1) has at least one minimizer. Further assume that SC holds at $x_{0}\in^{n}$. Let $F:^{n}\times^{m}\to^{n}$ be locally Lipschitz. Then, there exists $\delta_{0}>0$ such that the differential equation (18) has at least one solution $x:(-\delta_{0},\delta_{0})\to^{n}$ with initial condition $x(0)=x_{0}$.
@@
Note that since (1) has at least one minimizer for all $x\in^{n}$, the convexity assumptions on $f$ and $g$ imply that $u^{*}$ is a singleton for all $x\in^{n}$. Since SC holds at $x_{0}$, by [62, Prop. 5.39], since $g(x_{0},\cdot)$ is convex, MFCQ holds at $(x_{0},u^{*}(x_{0}))$. By [55, Theorem 5.3], this implies that $u^{*}$ is continuous at $x_{0}$. Since $g$ is continuous, there exists a neighborhood $\mathcal{V}_{x_{0}}$ of $x_{0}$ such that SC holds for all $y\in\mathcal{V}_{x_{0}}$. By the same argument, this implies that $u^{*}$ is continuous in $\mathcal{V}_{x_{0}}$. The result now follows by Peano’s existence theorem [69, Theorem 2.1]. ∎
Proposition 4.1 implies in particular that, under the assumptions of Proposition 3.2, the closed-loop system (18) has at least one solution in a neighborhood of $x_{0}$.
@@
Figure 2: The arrows depict the vector field (26). The dashed blue and orange curves depict the two solutions $y_{1}$ and $y_{2}$ starting from the origin, where the vector field is point-Lipschitz but not locally Lipschitz.
@@
If $\tilde{F}(x_{0})\neq 0$, by [70, Theorem 1.2.7], the differential equation has only one solution. If $\tilde{F}(x_{0})=0$, the result follows from Proposition 4.3. ∎
If Slater’s condition does not hold but the rest of assumptions of Proposition 3.2 hold (even with $f(x,\cdot)$ strongly convex for all $x\in^{n}$), Example 3.5 shows that $u^{*}$ can be discontinuous, in which case neither existence nor uniqueness of solutions is guaranteed. In the case where $f$ and $g$ are not differentiable with respect to the parameter, but the rest of assumptions of Proposition 3.2 hold (even with $f(x,\cdot)$ strongly convex for all $x\in^{n}$), Example 3.4 shows that $u^{*}$ is continuous but not necessarily point-Lipschitz. Therefore, in this case existence is guaranteed but uniqueness is not.
@@
(Existence of Filippov solutions for the closed-loop system): Suppose that $f$ and $g$ belong to $\mathcal{C}^{2}(^{n}\times^{m})$. Further assume that for any $x\in^{n}$, $f(x,\cdot)$ is strictly convex, $g(x,\cdot)$ is convex and (1) has at least one minimizer. Finally suppose that for every $x\in^{n}$, LCF holds at $x$ and $F:^{n}\times^{m}\to^{n}$ is locally Lipschitz. Then for any $x\in^{n}$, there exists $\delta_{x}>0$ such that (18) has at least one Filippov solution $y:(-\delta_{x},\delta_{x})\to^{n}$ with initial condition $y(0)=x$.
@@
Note that since (1) has at least one minimizer for all $x\in^{n}$, the convexity assumptions on $f$ and $g$ imply that $u^{*}$ is a singleton for all $x\in^{n}$. By Propositions 3.7 and 3.8, the assumptions ensure that $u^{*}$ is measurable and locally bounded. The result follows from [71, Theorem 7]. ∎
@@
Finally, we note that if $f$ is not strictly convex or $g$ is not convex, the optimizer $u^{*}$ is not guaranteed to be single-valued, which means that the usual notions of regularity of the controller and of solutions of the closed-loop system are not well defined. The third and fourth columns of Table 1 summarize the results presented in this section.